3) Find the solution to y -6y +8y=16 x(0)=0, x(0)=0 given that y.(x) = ce?* +cze** and y, (x) = 2.
Answer #1

- 6 8-3 +5° Ч®) = 09-0 Ч. et qe न girers that y(x) = Gent (х) = 2 - 1 (1) ** Ч yo get ge ta 90) = 5 + Сfa =0 aqe**+ 4G, е